I could really use some help with this one. I'm old and have forgotten how to do algebra!
5(x-1)=2x-26

Answer by 24HoursTutor.com(40)

5(x-1)=2x-26
You first need to open the brackets to get :
5x-5=2x-26
Now get all the x's on one side but remember when passing the equal sign the sign of the number changes
5x-2x=-26+5
3x=-21
x=-21/3
x=-7
Your answer is -7
The above question is solved by one of the experts from 24HoursTutor.com
We provide live 1:1 online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr